# Cosmic twins are likely rare: exploring pathways to hadron-quark stars with the same mass as neutron stars

### v1

@@ Acknowledgements.
We thank Thomas Baumgarte for useful discussions. We are grateful to Zachariah Etienne and Leo Werneck for releasing their work on the IllinoisGRMHD code which supports piecewise polytropic EOS, and to all the Einstein Toolkit developers and maintainers. We also thank Pedro Espino for sharing his EOS tables, and D. Alvarez-Castillo, D. Blanschke and A. Sedrakian for giving us permission to use the ACS-II equations of state developed in [34]. Our Python TOV solver for tabulated equations of state was developed starting from a code originally created by Nikolaos Stergioulas. Simulations were performed on the Puma cluster at the University of Arizona, and on ACCESS resources under allocation PHY190020. This work was supported in part by NSF grant PHY-2145421, and NASA grant 80NSSC24K0771 to the University of Arizona. G. B. was in part supported by the Frontera Fellowship by the Texas Advanced Computing Center (TACC). Frontera is funded by NSF grant No. OAC-1818253.
@@ Appendix A Extended EOS information
The EOSs used in this work consist of 8 segments and must satisfy the list of conditions in Sec. II. Empirically we have found that each $\Gamma_{i}$ controls specific characteristics of the TOV sequence shown in Fig. 1. $\Gamma_{1}$ primarily affects the maximum WD mass, $\Gamma_{2}$ alters the density and mass of the WD peak, $\Gamma_{i}$ for $i=\{3,4,5\}$ control the maximum NS mass, and $\Gamma_{6}$ changes the twin star mass range. Finally, $\Gamma_{i}$ for $i=\{7,8\}$ affects the maximum HS mass. However, it is important to emphasize that each of the free parameters affects more than just one characteristic.
The two EOSs listed in Table 1 exhibit only small differences, which are mostly noticeable at higher densities. Fig. 13 illustrates the pressure as a function of energy density $\epsilon$ for our EOSs. The shaded red region designates the density range where matter exists in the hadronic phase, and the blue region shows the density range where it exists purely in quark form. The narrower purple range, where the two regimes overlap ($i=6$ branch of the EOS), is where the phase transition occurs. In this regime, the original polytropic exponent was $\Gamma_{6}=0$, corresponding to a 1st-order phase transition, but we modified it to $\Gamma_{6}=0.2576$ to avoid a zero sound speed. This implies that, instead of a sharp boundary from the hadron to the quark phase, there exists a mixture of hadrons and quarks in this narrow region of our EOS. This modification makes the sound speed of this region $\simeq 15\,\%$ of the speed of light, which allows for stable numerical integration of the general relativistic hydrodynamic equations.
Figure 13: The relationship between the pressure and the energy density based on the EOSs constructed in Table 1. EOS I is represented by the solid line, while the dotted line denotes EOS II. The shaded red, blue, and purple areas correspond to the pure hadronic, pure quark, and the mixed quark-hadron phases for EOS I, respectively. The inset zooms in on the phase transition region.
@@ Appendix B Cooling formalism
where $\alpha$, $\beta^{i}$, and $\gamma_{ij}$ are the lapse function, shift vector, and spatial metric ($i$ and $j$ are spatial indices). For a perfect fluid with 4-velocity $u^{\mu}$, the stress-energy tensor is written as
@@
$\begin{gathered}\partial_{t}\rho_{\ast}+\partial_{j}(\rho_{\ast}v^{j})=0\;,\\[10.00002pt] \partial_{t}\tilde{\tau}+\partial_{i}(\alpha^{2}\sqrt{\gamma}\;T^{0i}-\rho_{\ast}v^{i})=s\;,\\[10.00002pt] \partial_{t}\tilde{S}_{i}+\partial_{j}(\alpha\sqrt{\gamma}\;T^{j}_{\;\;i})=\frac{1}{2}\alpha\sqrt{\gamma}\;T^{\mu\nu}\;\partial_{i}g_{\mu\nu}\;,\\[10.00002pt] \end{gathered}$ (13)
In these equations, $v^{i}=u^{i}/u^{0}$ is the coordinate three velocity, and the conservative variables are defined as
@@
where $n^{\mu}=\frac{1}{\alpha}(1,-\beta^{i})$ is the timelike unit vector normal to $t=$ constant slices.
@@
The set of Eqs. (LABEL:GRMHD) also requires an EOS to close. The general form of the EOS that is assumed to solve this set of equations is given by Eq. (9[143], where $P_{cold}$ stands for the pressure of the zero-temperature matter, and has the polytropic form introduced in Sec. II. Any additional pressure due to heating, e.g., by shock heating, is encapsulated in $P_{th}$. This term is considered as a $\Gamma$-law,
@@
In the presence of radiation, the conservation of energy-momentum must be modified to incorporate the radiation stress-energy tensor $R^{\mu\nu}$,
@@
The dynamics of the radiation stress-energy tensor is described by $\nabla_{\mu}R^{\mu\nu}=-G^{\nu}$, where $G^{\mu}$ is the radiation 4-force density. Assuming local and isotropic (in a frame comoving with the fluid) cooling, the resulting energy and momentum equations are then given by [124]
@@ Appendix C Cooling tests
Figure 14: The value of $|e_{th}|$ is shown in 2-dimensional snapshots of three simulations involving a stable NS governed by a $\Gamma=2$ polytropic EOS with constant $k=100\,M_{\odot}^{2}$, and $\rho_{c}=0.001\,M_{\odot}^{-2}$ The evolution of each simulation is illustrated in a separate row. The first row describes an NS without any initial perturbation in pressure and without cooling. The second row corresponds to the case with a small pressure perturbation in the initial data and no cooling. The bottom row has the same initial data as the second row but cooling is active. The plot demonstrates that cooling is effective at removing excess heat in the bulk of the stellar matter. The stellar surface discontinuity generates heat at every timestep and would require more aggressive cooling to completely cool. However, the amount of mass in that hotter region is negligible compared to that of the cold bulk of the star.
Fig. 14 shows the result of these simulations for the absolute value of $e_{th}$ at four different times. The three simulations are presented in the three rows, respectively. Even in the first simulation, where there is no additional thermal pressure in the initial data, small numerical errors result in the generation of some heat that grows over time. These numerical errors can also deplete heat, leading to negative $e_{th}$ at some points. This is why we show the absolute value of this quantity. The initial data for this simulation contains no initial heat, as there is no heat source of $P_{th}$ at the beginning. In the second simulation, the perturbation $\xi_{p}=0.01$ results in heat in the initial data. This excess heat can accumulate alongside the contribution from the numerical errors over time, resulting in a relatively hotter configuration. The third simulation starts with the same amount of initial heat as the second simulation, but cooling removes the extra heat from the star. As can be seen in the bottom row of plots in Fig. 14, the initial heat and a significant portion of the heat arising from numerical errors are effectively removed after a few dynamical times.
@@
Figure 17: The evolution of the quantity $e_{th}/e$ at the center of an NS during the early times. The green curve represents the simulation results with $\xi_{p}=0.01$ and $\tau_{c}=3.16\,t_{dyn}$. The fitted model described by Eq. (25) is shown by dots, with the damping timescale set to $\tau_{damp}=3.20\,t_{dyn}$.
@@ Appendix D Stable stars
For these simulations, the grid structure consists of 8 and 7 refinement levels, respectively. In both configurations, the finest resolution is set to 100 points across the initial radius of the stars. The half-side length of the coarsest box is $134.4R$ for the stable NS simulation and $67.2R$ for the stable HS simulation.
Fig. 18 shows the relative deviation of the central density over time, with respect to its initial value. This figure indicates that over $8\,t_{dyn}$, the stable NS exhibits density fluctuations of $\sim 0.01\%$ at maximum, while the largest amplitude of fluctuations for the HS is $\sim 0.1\%$. Due to the transition to the more complex part of the EOS, the density evolution of the stable HS follows a less predictable pattern.
@@ Appendix E Constraints
The absolute value of the constraint violation averaged over volume as a function of time is shown in Fig. 19 for both a stable NS (solid curves) and a collapsing WD with cooling activated from the beginning (dotted curves). The constraint violations are denoted by $\Delta_{k}$. Red corresponds to the violation of the Hamiltonian constraint $\mathcal{H}$, while the momentum constraints violations, $\mathcal{M}_{j}$ with $j=\{x,y,z\}$, correspond to orange, green, and blue. $\Delta_{k}$ should remain close to zero at all timesteps. As the star undergoes collapse, the computational errors increase, leading to larger values of $\Delta_{k}$. Since $\Delta_{k}$ is a dimensionful quantity, we also show $\Delta_{k}$ for the stable NS simulation discussed in App. D for comparison. The plot demonstrates that the constraints remain well satisfied throughout the collapse. The maximum value of $\Delta_{k}$ remains small even in the simulations with more aggressive initial perturbations presented in this work, and it similarly occurs during the late stages of the collapse. The worst case in our simulations involved the HS with mass loss, where constraint violations are still $\Delta_{k}\lesssim 10^{-9}\,M_{\odot}^{-2}$.

### v2

@@ Acknowledgements.
We thank Thomas Baumgarte for useful discussions. We are grateful to Zachariah Etienne and Leo Werneck for releasing their work on the IllinoisGRMHD code which supports piecewise polytropic EOS, and to all the Einstein Toolkit developers and maintainers. We also thank Pedro Espino for sharing his EOS tables, and D. Alvarez-Castillo, D. Blaschke and A. Sedrakian for giving us permission to use the ACS-II equations of state developed in [34]. Our Python TOV solver for tabulated equations of state was developed starting from a code originally created by Nikolaos Stergioulas. Simulations were performed on the Puma cluster at the University of Arizona, and on ACCESS resources under Allocation No. PHY190020. This work was supported in part by NSF Grant No. PHY-2145421 and NASA Grant No. 80NSSC24K0771 to the University of Arizona. G. B. was in part supported by the Frontera Fellowship by the Texas Advanced Computing Center (TACC). Frontera is funded by NSF Grant No. OAC-1818253.
@@ Appendix A Extended EOS information
The EOSs used in this work consist of eight segments and must satisfy the list of conditions in Sec. II. Empirically we have found that each $\Gamma_{i}$ controls specific characteristics of the TOV sequence shown in Fig. 1. $\Gamma_{1}$ primarily affects the maximum WD mass, $\Gamma_{2}$ alters the density and mass of the WD peak, $\Gamma_{i}$ for $i=\{3,4,5\}$ controls the maximum NS mass, and $\Gamma_{6}$ changes the twin star mass range. Finally, $\Gamma_{i}$ for $i=\{7,8\}$ affects the maximum HS mass. However, it is important to emphasize that each of the free parameters affects more than just one characteristic.
The two EOSs listed in Table 1 exhibit only small differences, which are mostly noticeable at higher densities. Fig. 13 illustrates the pressure as a function of energy density $\epsilon$ for our EOSs. The shaded red region designates the density range where matter exists in the hadronic phase, and the blue region shows the density range where it exists purely in quark form. The narrower purple range, where the two regimes overlap ($i=6$ branch of the EOS), is where the phase transition occurs. In this regime, the original polytropic exponent was $\Gamma_{6}=0$, corresponding to a first-order phase transition, but we modified it to $\Gamma_{6}=0.2576$ to avoid a zero sound speed. This implies that, instead of a sharp boundary from the hadron to the quark phase, there exists a mixture of hadrons and quarks in this narrow region of our EOS. This modification makes the sound speed of this region $\simeq 15\,\%$ of the speed of light, which allows for stable numerical integration of the general relativistic hydrodynamic equations.
Figure 13: The relationship between the pressure and the energy density based on the EOSs constructed in Table 1. EOS I is represented by the solid line, while the dotted line denotes EOS II. The shaded red, blue, and purple areas correspond to the pure hadronic, pure quark, and the mixed quark-hadron phases for EOS I, respectively. The inset enlarges the phase transition region.
@@ Appendix B Cooling formalism
where $\alpha$, $\beta^{i}$, and $\gamma_{ij}$ are the lapse function, shift vector, and spatial metric, respectively ($i$ and $j$ are spatial indices). For a perfect fluid with 4-velocity $u^{\mu}$, the stress-energy tensor is written as
@@
$\begin{gathered}\partial_{t}\rho_{\ast}+\partial_{j}(\rho_{\ast}v^{j})=0\;,\\[10.00002pt] \partial_{t}\tilde{\tau}+\partial_{i}(\alpha^{2}\sqrt{\gamma}\;T^{0i}-\rho_{\ast}v^{i})=s\;,\\[10.00002pt] \partial_{t}\tilde{S}_{i}+\partial_{j}(\alpha\sqrt{\gamma}\;T^{j}_{\;\;i})=\frac{1}{2}\alpha\sqrt{\gamma}\;T^{\mu\nu}\;\partial_{i}g_{\mu\nu}\;.\\[10.00002pt] \end{gathered}$ (13)
In these equations, $v^{i}=u^{i}/u^{0}$ is the coordinate three-velocity, and the conservative variables are defined as
@@
where $n^{\mu}=\frac{1}{\alpha}(1,-\beta^{i})$ is the timelike unit vector normal to $t=$ const slices.
@@
The set of Eqs. (LABEL:GRMHD) also requires an EOS to close. The general form of the EOS that is assumed to solve this set of equations is given by Eq. (9[150], where $P_{cold}$ stands for the pressure of the zero-temperature matter, and has the polytropic form introduced in Sec. II. Any additional pressure due to heating, e.g., by shock heating, is encapsulated in $P_{th}$. This term is considered as a $\Gamma$-law:
@@
In the presence of radiation, the conservation of energy-momentum must be modified to incorporate the radiation stress-energy tensor $R^{\mu\nu}$:
@@
The dynamics of the radiation stress-energy tensor is described by $\nabla_{\mu}R^{\mu\nu}=-G^{\nu}$, where $G^{\mu}$ is the radiation four-force density. Assuming local and isotropic (in a frame comoving with the fluid) cooling, the resulting energy and momentum equations are then given by [124]
@@ Appendix C Cooling tests
Figure 14: The value of $|e_{th}|$ is shown in two-dimensional snapshots of three simulations involving a stable NS governed by a $\Gamma=2$ polytropic EOS with constant $k=100\,M_{\odot}^{2}$, and $\rho_{c}=0.001\,M_{\odot}^{-2}$ The evolution of each simulation is illustrated in a separate row. The first row describes a NS without any initial perturbation in pressure and without cooling. The second row corresponds to the case with a small pressure perturbation in the initial data and no cooling. The bottom row has the same initial data as the second row but cooling is active. The plot demonstrates that cooling is effective at removing excess heat in the bulk of the stellar matter. The stellar surface discontinuity generates heat at every time step and would require more aggressive cooling to completely cool. However, the amount of mass in that hotter region is negligible compared to that of the cold bulk of the star.
Fig. 14 shows the result of these simulations for the absolute value of $e_{th}$ at four different times. The three simulations are presented in the three rows, respectively. Even in the first simulation, where there is no additional thermal pressure in the initial data, small numerical errors result in the generation of some heat that grows over time. These numerical errors can also deplete heat, leading to negative $e_{th}$ at some points. This is why we show the absolute value of this quantity. The initial data for this simulation contain no initial heat, as there is no heat source of $P_{th}$ at the beginning. In the second simulation, the perturbation $\xi_{p}=0.01$ results in heat in the initial data. This excess heat can accumulate alongside the contribution from the numerical errors over time, resulting in a relatively hotter configuration. The third simulation starts with the same amount of initial heat as the second simulation, but cooling removes the extra heat from the star. As can be seen in the bottom row of plots in Fig. 14, the initial heat and a significant portion of the heat arising from numerical errors are effectively removed after a few dynamical times.
@@
Figure 17: The evolution of the quantity $e_{th}/e$ at the center of a NS during the early times. The green curve represents the simulation results with $\xi_{p}=0.01$ and $\tau_{c}=3.16\,t_{dyn}$. The fitted model described by Eq. (25) is shown by dots, with the damping timescale set to $\tau_{damp}=3.20\,t_{dyn}$.
@@ Appendix D Stable stars
For these simulations, the grid structure consists of eight and seven refinement levels, respectively. In both configurations, the finest resolution is set to 100 points across the initial radius of the stars. The half-side length of the coarsest box is $134.4R$ for the stable NS simulation and $67.2R$ for the stable HS simulation.
Fig. 18 shows the relative deviation of the central density over time, with respect to its initial value. This figure indicates that over $8\,t_{dyn}$, the stable NS exhibits density fluctuations of $\sim 0.01\%$ at maximum, while the largest amplitude of fluctuations for the HS is $\sim 0.1\%$. Because of the transition to the more complex part of the EOS, the density evolution of the stable HS follows a less predictable pattern.
@@ Appendix E Constraints
The absolute value of the constraint violation averaged over volume as a function of time is shown in Fig. 19 for both a stable NS (solid curves) and a collapsing WD with cooling activated from the beginning (dotted curves). The constraint violations are denoted by $\Delta_{k}$. Red corresponds to the violation of the Hamiltonian constraint $\mathcal{H}$, while the momentum constraint violations, $\mathcal{M}_{j}$ with $j=\{x,y,z\}$, correspond to orange, green, and blue. $\Delta_{k}$ should remain close to zero at all time steps. As the star undergoes collapse, the computational errors increase, leading to larger values of $\Delta_{k}$. Since $\Delta_{k}$ is a dimensionful quantity, we also show $\Delta_{k}$ for the stable NS simulation discussed in App. D for comparison. The plot demonstrates that the constraints remain well satisfied throughout the collapse. The maximum value of $\Delta_{k}$ remains small even in the simulations with more aggressive initial perturbations presented in this work, and it similarly occurs during the late stages of the collapse. The worst case in our simulations involved the HS with mass loss, where constraint violations are still $\Delta_{k}\lesssim 10^{-9}\,M_{\odot}^{-2}$.